# A contextual advantage for conclusive exclusion: repurposing the Pusey-Barrett-Rudolph construction

### v1

@@ Abstract
Introduction— Conclusive exclusion refers to a quantum state discrimination task where one is given a set $\{\rho_{k}\}_{k=1}^{n}$ of quantum states, and one must find a measurement $\{E_{k}\}_{k=1}^{n}$ (i.e., with a number of outcomes equal to the number of states in the set) such that obtaining outcome $E_{k}$ allows one to conclude with certainty that state $\rho_{k}$ was not prepared. The success probability for conclusive exclusion can be quantified by
@@
where $\mathbb{P}(E_{k}|\rho_{k})$ is the probability of obtaining outcome $E_{k}$ when the state is $\rho_{k}$. Conclusive exclusion occurs when this quantity is one.
@@
The task of conclusive exclusion was first studied by Caves, Fuchs and Schack in an article on quantum state compatibility (under the name “post-Peierls compatibility”) [CFS2002]. It subsequently became a subject of greater interest when it was used by Pusey, Barrett and Rudolph (PBR) as part of an argument about the interpretation of the quantum state [puseyReality2012]. The term “conclusive exclusion” was introduced in Ref. [PerryJainOppenheim], where a study was undertaken of what conditions a set of states must satisfy in order for there to be a possibility of conclusive exclusion. Leifer’s review [leifer2014quantum] of $\psi$-ontology theorems highlighted the significance of conclusive exclusion (under the name ‘antidistinguishability”).
@@
We undertake a study of what aspects of conclusive exclusion resist classical explanation. The notion of classicality of experimental statistics we adopt here is realizability by a generalized-noncontextual (henceforth, simply ‘noncontextual’) ontological model [harriganEinstein2010, Spekkens2005], whose motivations are given in Refs. [spekkens2019, Spekkens2005, schmid2021guiding, Schmid2021].
@@
The causal structure of the experiment considered by PBR involves two systems, $A$ and $B$, that are prepared independently and measured jointly. We consider a slight generalization of this causal structure, depicted in  Fig. 1, which we refer to as the $\rm PBR$$^{+}$ scenario’ [^1] Like the original PBR paper, we consider two quantum systems prepared independently and measured jointly; however, we conceptualize the different preparations and measurements via multi-sources and multi-meters.. We show that the quantum prediction for how well one can achieve conclusive exclusion within the $\rm PBR$$^{+}$ scenario resists classical explanation. We do this by showing that such conclusive exclusion implies a novel proof of the impossibility of a noncontextual ontological model. We moreover derive robust noncontextuality inequalities bounding how well one can achieve this task, and we show that quantum theory can violate these bounds. These results help tease out what is genuinely nonclassical about the phenomenology of quantum state discrimination, complementing prior works showing quantum-over-classical advantages for minimum error and unambiguous state discrimination [schmidContextual2018, statedisc2, statedisc3, statedisc4, statedisc5, statedisc6, flatt2025unifyingcontextualadvantagesstate]. They furthermore contribute to the broader project of identifying the precise boundary between classical and quantum for any given phenomenology, as has been carried out in the contexts of computation [Schmid2022Stabilizer, shahandeh2021quantum], interference [Catani2023whyinterference, catani2022reply, catani2023aspects], compatibility [selby2023incompatibility, selby2023accessible, PhysRevResearch.2.013011], uncertainty relations [catani2022nonclassical], metrology [contextmetrology], thermodynamics [contextmetrology, comar2024contextuality], weak values [AWV, KLP19], coherence [rossi2023contextuality, Wagner2024coherence], quantum Darwinism [baldijao2021noncontextuality], information processing and communication [POM, RAC, RAC2, Saha_2019, Yadavalli2020, PhysRevLett.119.220402, fonseca2024robustness], cloning [cloningcontext], broadcasting [jokinen2024nobroadcasting], as well as Bell scenarios [Wright2023invertible, schmid2020unscrambling] and Kochen-Specker scenarios [operationalks, kunjwal2018from, Kunjwal16, Kunjwal19, Kunjwal20, specker, Gonda2018almostquantum].
@@
We take the two systems in the $\rm PBR$$^{+}$ scenario to be two qubits and denote them $A$ and $B$. We now define the construction, beginning with a particular choice of product states on the composite $AB$. The basic construction presumes that each qubit has a state chosen from the set $\Omega_{0+}\coloneqq\{|0\rangle,|+\rangle\},$ so that the possible product states for the two qubits form the set denoted $\mathcal{P}_{0+}\coloneqq\{\rho_{k|0+}\}_{k}$ with (nonorthogonal) elements
@@
$\rho_{k|0-}\coloneqq(Z\otimes Z)\rho_{k|0-},$ (4a)
$\rho_{k|1+}\coloneqq(X\otimes X)\rho_{k|1+},$ (4b)
$\rho_{k|1-}\coloneqq(Y\otimes Y)\rho_{k|1-}.$ (4c)
@@
$E_{k|0-}\coloneqq(Z\otimes Z)E_{k|0-},$ (5a)
$E_{k|1+}\coloneqq(X\otimes X)E_{k|1+},$ (5b)
$E_{k|1-}\coloneqq(Y\otimes Y)E_{k|1-}.$ (5c)
@@
The 16 possible product states on $AB$ that this pair of multi-sources can generate, as a function of the different possible values of $(S_{A},X_{A};S_{B},X_{B})$, are listed in Fig. 2. All of the states in our four conclusive exclusion protocols can be so generated, and so appear in the table. (Note that four of the product states in the table do not appear in any of our conclusive exclusion protocols, while the four along the diagonal of the table each appear in two of the conclusive exclusion protocol.) The states appearing in $\mathcal{P}_{0+}$ are highlighted in yellow, those in $\mathcal{P}_{0-}$ in red, those in $\mathcal{P}_{1+}$ in green, and those in $\mathcal{P}_{1-}$ are highlighted in blue.
@@
${\rm CE}_{T}=1-\frac{1}{4}\sum_{Y}\mathbb{P}(E_{Y|T}|\tilde{\rho}_{S_{A}(Y,T),X_{A}(Y,T),S_{B}(Y,T),Y_{B}(Y,T)}),$ (7)
where $S_{A}(Y,T)$ indicates the value of $S_{A}$ as a function of the tuple $(Y,T)$, and similarly for $X_{A}(Y,T),S_{B}(Y,T),Y_{B}(Y,T)$, such that $\tilde{\rho}_{S_{A}(Y,T),X_{A}(Y,T),S_{B}(Y,T),Y_{B}(Y,T)}=\rho_{Y|T}$. (The explicit form of these functions can be identified by checking how the $\rho_{Y|T}$s defined in Eqs. 2 and 4 correspond to states generated by $(S_{A},X_{A},S_{B},X_{B})$ as shown in Fig. 2.)
@@
Connection to the bilocality scenario— Assumptions about the causal structure that holds among a number of variables can imply constraints on the correlations that can be realized on these variables, with the paradigm example being how the Bell causal structure (with a classical latent variable as a common cause) implies that the distribution over outcome variables given setting variables should satisfy Bell inequalities. Generally, inequality constraints on the realizable correlations when the latent nodes are assumed to be classical in a given causal structure are termed classical causal compatibility inequalities for that causal structure [Wolfe2016inflation] (see Refs. [pearl2009causality, fritz2012beyond, Himbeeck2018instrumental] for examples). In Appendix C, we show that the noncontextuality inequality of Eq. 12 for the $\rm PBR$$^{+}$ scenario (Fig. 1) arises as a classical causal compatibility inequality within the causal structure known as the bilocality scenario [bilocality], depicted in Fig. 4. (Recall that from Eq. (7), this inequality can be expressed in terms of the conditional probability $\mathbb{P}{(X_{A}X_{B}Y|S_{A}S_{B}T)}$ that are native to the bilocality scenario.) That is, if one is certain that the causal structure of one’s experiment is as shown in Fig. 4, then one does not even need to appeal to noncontextuality to derive the bound in Eq. (12)—rather, the bound follows from the causal structure itself.
@@
One way to obtain a nontrivial classical causal compatibility inequality (i.e., one that can be violated by quantum theory) in the bilocality scenario is to piggyback on Bell’s theorem: by postselecting on the joint measurement’s outcome, one can achieve entanglement swapping, effectively reducing the scenario to a standard Bell scenario with entanglement between $A^{\prime}$ and $B^{\prime}$. It follows that one can achieve every variety of proof that is achievable in the standard Bell scenario, including possibilistic proofs (the paradigm example of which is Hardy’s proof [Hardy1993paradox]) wherein the quantum-classical gap is manifest already at the level of the support of the distribution, i.e., a specification of which outcomes are possible and which are impossible, without regard to the values of the nonzero probabilities. However, the question of interest has always been whether there are novel proofs of quantum-classical gaps in the bilocality scenario, i.e., proofs which do not piggyback on Bell’s theorem through the entanglement swapping protocol [bilocality, BilocalCorrelations, ciudadalañón2024escapingshadowbellstheorem]. The quantum-classical gap we have manifested here is of this variety. Furthermore, to our knowledge, it is the first example of a novel proof of nonclassicality in the bilocality scenario that is possibilistic.
@@
Acknowledgements— DS, RWS, and YY were supported by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. YY was also supported by the Natural Sciences and Engineering Research Council of Canada (Grant No. RGPIN-2024-04419).
@@ Appendix A Overlapping ontic supports
As proven in Ref. [schmidContextual2018], $\Lambda_{0}\neq\Lambda_{+}$, thus, there exists $\lambda^{\ast}\in\Lambda_{0}$ such that $\lambda^{\ast}\notin\Lambda{+}$, i.e., $\mu_{0}(\lambda^{\ast})>0$ while $\mu_{+}(\lambda^{\ast})=0$. Then, together with the fact that $\mu_{1}(\lambda)$ is nonnegative, Eq. 16 implies that $\mu_{-}(\lambda^{\ast})>0$. Thus, $\lambda^{\ast}$ is in the ontic support for both $\mu_{0}$ and $\mu_{-}$, implying that $\Omega_{0-}$ has nontrivial ontic overlap.
@@ Appendix B Proof of the noncontextuality inequality
Here we derive the noncontextuality inequality on CE, namely, Eq. 12, and show that it is tight, meaning that the maximal value of CE that can be achieved for any theory that admits of a noncontextual ontological model is 3. Such a proof does not assume the correctness of quantum theory, so we first translate the scenario described in the main text to one for a general GPT.
@@
$\frac{1}{2}\vec{s}_{0}+\frac{1}{2}\vec{s}_{1}=\frac{1}{2}\vec{s}_{+}+\frac{1}{2}\vec{s}_{-},$ (13)
@@
${\rm CE}_{0+}\coloneqq 1-\frac{1}{4}\sum^{4}_{k=1}\mathbb{P}(\vec{e}_{Y|T}|\vec{s}_{Y|T})=1-\frac{1}{4}\sum^{4}_{k=1}\vec{e}_{Y|T}\cdot\vec{s}_{Y|T}.$ (14)
Summing the probabilities of conclusive exclusion for the four tasks, we have the quantity for which we will provide a noncontextual bound, namely,
@@
$\frac{1}{2}\mu_{0}(\lambda)+\frac{1}{2}\mu_{1}(\lambda)=\frac{1}{2}\mu_{+}(\lambda)+\frac{1}{2}\mu_{-}(\lambda)$ (16)
The ontic state space of a bipartite system is simply the Cartesian product, namely, $\Lambda_{A}\times\Lambda_{B}=\Lambda\times\Lambda$, and the ontic representation of a product state is the product of the ontic representations of the unipartite state in the product. We denote the response function representing the measurement effect $\vec{e}_{Y|T}$ as $\xi_{Y|T}(\lambda_{A},\lambda_{B})$ to, and denote the probability distribution representing the GPT state on $AB$, $\vec{s}_{Y|T}$, as $\mu_{Y|T}(\lambda_{A},\lambda_{B})$, such that
@@
${\rm CE}_{0+}\leq$ $1-\int_{\Lambda_{0}\cap\Lambda_{+}}\int_{\Lambda_{0}\cap\Lambda_{+}}[\xi_{00|0+}(\lambda_{A},\lambda_{B})\mu_{0}(\lambda_{A})\mu_{0}(\lambda_{B})+\xi_{0+|0+}(\lambda_{A},\lambda_{B})\mu_{0}(\lambda_{A})\mu_{+}(\lambda_{B})$
$+\xi_{+0|0+}(\lambda_{A},\lambda_{B})\mu_{+}(\lambda_{A})\mu_{0}(\lambda_{B})+\xi_{++|0+}(\lambda_{A},\lambda_{B})\mu_{+}(\lambda_{A})\mu_{+}(\lambda_{B})]d\lambda_{A}d\lambda_{B}$
$\leq$ $1-\frac{1}{4}\int_{\Lambda_{0}\cap\Lambda_{+}}\int_{\Lambda_{0}\cap\Lambda_{+}}\sum^{4}_{Y=1}\xi_{Y|0+}(\lambda_{A},\lambda_{B})\min\{\mu_{0}(\lambda_{A}),\mu_{+}(\lambda_{A})\}\min\{\mu_{0}(\lambda_{B}),\mu_{+}(\lambda_{B})\}d\lambda_{A}d\lambda_{B}$ (20)
$=$ $1-\frac{1}{4}\left[\int_{\Lambda_{0}\cap\Lambda_{+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2}.$
@@
${\rm CE}_{0-}\leq 1-\frac{1}{4}\left[\int_{\Lambda_{0}\cap\Lambda_{-}}\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2},$
${\rm CE}_{1+}\leq 1-\frac{1}{4}\left[\int_{\Lambda_{1}\cap\Lambda_{+}}\min\{\mu_{1}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2},$
${\rm CE}_{0-}\leq 1-\frac{1}{4}\left[\int_{\Lambda_{1}\cap\Lambda_{-}}\min\{\mu_{1}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2}.$
@@
${\rm CE}\leq 1-\frac{1}{4}\left\{\left[\int_{\Lambda_{0}\cap\Lambda_{+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2}+\left[\int_{\Lambda_{0}\cap\Lambda_{-}}\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2}\right$. (22)
$\left.+\left[\int_{\Lambda_{1}\cap\Lambda_{+}}\min\{\mu_{1}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2}+\left[\int_{\Lambda_{1}\cap\Lambda_{-}}\min\{\mu_{1}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2}\right\}.$
@@
$\beta_{0+}\coloneqq\int_{\Lambda_{0}\cap\Lambda_{+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda,\quad\beta_{0-}\coloneqq\int_{\Lambda_{0}\cap\Lambda_{-}}\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}d\lambda,$
$\beta_{1+}\coloneqq\int_{\Lambda_{1}\cap\Lambda_{+}}\min\{\mu_{1}(\lambda),\mu_{+}(\lambda)\}d\lambda,\quad\beta_{1-}\coloneqq\int_{\Lambda_{1}\cap\Lambda_{-}}\min\{\mu_{1}(\lambda),\mu_{-}(\lambda)\}d\lambda.$
@@
$\big(\beta_{0+}^{2}+\beta_{0-}^{2}+\beta_{1+}^{2}+\beta_{1-}^{2}\big)\big(1^{2}+1^{2}+1^{2}+1^{2}\big)\geq\big(\beta_{0+}+\beta_{0-}+\beta_{1+}+\beta_{1-}\big)^{2}=4.$ (25)
@@
${\rm CE}\leq 1-\frac{1}{4}\big(\beta_{0+}+\beta_{0-}+\beta_{1+}+\beta_{1-}\big)^{2}.$ (26)
@@
$\beta_{0+}=$ $\int_{\Lambda_{0+}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{+}\}d\lambda.$
@@
$\Lambda_{1+}\coloneqq\Lambda_{1}\cap\Lambda_{+}\setminus\Lambda_{01+-}\cup\Lambda_{01+}\cup\Lambda_{1+-},$ (30)
$\Lambda_{1-}\coloneqq\Lambda_{1}\cap\Lambda_{-}\setminus\Lambda_{01+-}\cup\Lambda_{01-}\cup\Lambda_{1+-},$ (31)

### v2

@@ Abstract
Introduction— Conclusive exclusion refers to a quantum state discrimination task where one is given a set $\{\rho_{k}\}_{k=1}^{n}$ of quantum states, and one must find a measurement $\{E_{k}\}_{k=1}^{n}$ (i.e., with a number of outcomes equal to the number of states in the set) such that obtaining outcome $E_{k}$ allows one to conclude with certainty that state $\rho_{k}$ was not prepared. The measure of success for conclusive exclusion can be quantified by
@@
where $\mathbb{P}(E_{k}|\rho_{k})$ is the probability of obtaining outcome $E_{k}$ when the state is $\rho_{k}$. Perfect conclusive exclusion occurs when this quantity is one.
@@
The task of conclusive exclusion was first studied by Caves, Fuchs, and Schack in an article on quantum state compatibility (under the name “post-Peierls compatibility”) [CFS2002]. It subsequently became a subject of greater interest when it was used by Pusey, Barrett, and Rudolph (PBR) as part of an argument about the interpretation of the quantum state [puseyReality2012]. The term “conclusive exclusion” was introduced in Ref. [PerryJainOppenheim], where a study was undertaken of what conditions a set of states must satisfy in order for there to be a possibility of conclusive exclusion. Leifer’s review [leifer2014quantum] of $\psi$-ontology theorems highlighted the significance of conclusive exclusion (under the name “antidistinguishability”).
@@
We undertake a study of what aspects of conclusive exclusion resist classical explanation. The notion of classicality of experimental statistics we adopt here is realizability by a generalized-noncontextual (henceforth, simply “noncontextual”) ontological model [harriganEinstein2010, Spekkens2005], whose motivations are given in Refs. [spekkens2019, Spekkens2005, schmid2021guiding, Schmid2021].
@@
The causal structure of the experiment considered by PBR involves two systems, $A$ and $B$, that are prepared independently and measured jointly. We consider a slight generalization of this causal structure, depicted in Fig. 1, which we refer to as the $\rm PBR$$^{+}$ scenario” [^1] Like the original PBR paper, we consider two quantum systems prepared independently and measured jointly; however, we conceptualize the different preparations and measurements via multi-sources and multi-meters.. We show that the quantum prediction for how well one can achieve conclusive exclusion within the $\rm PBR$$^{+}$ scenario resists classical explanation. We do this by showing that such conclusive exclusion implies a novel proof of the impossibility of a noncontextual ontological model. We moreover derive robust noncontextuality inequalities bounding how well one can achieve this task, and we show that quantum theory can violate these bounds. These results help tease out what is genuinely nonclassical about the phenomenology of quantum state discrimination, complementing prior works showing quantum-over-classical advantages for minimum error and unambiguous state discrimination [schmidContextual2018, statedisc2, statedisc3, statedisc4, statedisc5, statedisc6, flatt2025unifyingcontextualadvantagesstate]. They furthermore contribute to the broader project of identifying the precise boundary between classical and quantum for any given phenomenology, as has been carried out in the contexts of computation [Schmid2022Stabilizer, shahandeh2021quantum], interference [Catani2023whyinterference, catani2022reply, catani2023aspects], compatibility [selby2023incompatibility, selby2023accessible, PhysRevResearch.2.013011], uncertainty relations [catani2022nonclassical], metrology [contextmetrology], thermodynamics [contextmetrology, comar2024contextuality], weak values [AWV, KLP19], coherence [rossi2023contextuality, Wagner2024coherence], quantum Darwinism [baldijao2021noncontextuality], information processing and communication [POM, RAC, RAC2, Saha_2019, Yadavalli2020, PhysRevLett.119.220402, fonseca2024robustness], cloning [cloningcontext], broadcasting [jokinen2024nobroadcasting], as well as Bell scenarios [Wright2023invertible, schmid2020unscrambling] and Kochen-Specker scenarios [operationalks, kunjwal2018from, Kunjwal16, Kunjwal19, Kunjwal20, specker, Gonda2018almostquantum].
@@
We take the two systems in the $\rm PBR$$^{+}$ scenario to be two qubits and denote them $A$ and $B$. We now define the construction, beginning with a particular choice of product states on the composite $AB$. The basic construction presumes that each qubit has a state chosen from the set $\Omega_{0+}\coloneqq\{|0\rangle,|+\rangle\}$ so that the possible product states for the two qubits form the set denoted $\mathcal{P}_{0+}\coloneqq\{\rho_{k|0+}\}_{k=1}^{4}$ with (nonorthogonal) elements
@@
$\rho_{k|0-}\coloneqq(Z\otimes Z)\rho_{k|0+}(Z\otimes Z),$ (4a)
$\rho_{k|1+}\coloneqq(X\otimes X)\rho_{k|0+}(X\otimes X),$ (4b)
$\rho_{k|1-}\coloneqq(Y\otimes Y)\rho_{k|0+}(Y\otimes Y).$ (4c)
@@
$E_{k|0-}\coloneqq(Z\otimes Z)E_{k|0+}(Z\otimes Z),$ (5a)
$E_{k|1+}\coloneqq(X\otimes X)E_{k|0+}(X\otimes X),$ (5b)
$E_{k|1-}\coloneqq(Y\otimes Y)E_{k|0+}(Y\otimes Y).$ (5c)
@@
The 16 possible product states on $AB$ that this pair of multi-sources can generate, as a function of the different possible values of $(S_{A},X_{A};S_{B},X_{B})$, are listed in Fig. 2. All of the states in our four conclusive exclusion protocols can be so generated, and so appear in the table. (Note that four of the product states in the table do not appear in any of our conclusive exclusion protocols, while the four along the diagonal of the table each appear in two of the conclusive exclusion protocols.) The states appearing in $\mathcal{P}_{0+}$ are highlighted in yellow, those in $\mathcal{P}_{0-}$ in red, those in $\mathcal{P}_{1+}$ in green, and those in $\mathcal{P}_{1-}$ are highlighted in blue.
@@
${\rm CE}_{T}=1-\frac{1}{4}\sum_{Y}\mathbb{P}(E_{Y|T}|\tilde{\rho}_{S_{A}(Y,T),X_{A}(Y,T),S_{B}(Y,T),X_{B}(Y,T)}),$ (7)
where $S_{A}(Y,T)$ indicates the value of $S_{A}$ as a function of the tuple $(Y,T)$, and similarly for $X_{A}(Y,T)$, $S_{B}(Y,T)$ and $X_{B}(Y,T)$, such that $\tilde{\rho}_{S_{A}(Y,T),X_{A}(Y,T),S_{B}(Y,T),X_{B}(Y,T)}=\rho_{Y|T}$. (The explicit form of these functions can be identified by checking how the $\rho_{Y|T}$s defined in Eqs. 2 and 4 correspond to states generated by $(S_{A},X_{A},S_{B},X_{B})$ as shown in Fig. 2.)
@@
Connection to the bilocality scenario— Assumptions about the causal structure that holds among a number of variables can imply constraints on the correlations that can be realized on these variables, with the paradigm example being how the Bell causal structure (with a classical latent variable as a common cause) implies that the distribution over outcome variables given setting variables should satisfy Bell inequalities. Generally, inequality constraints on the realizable correlations when the latent nodes are assumed to be classical in a given causal structure are termed classical causal compatibility inequalities for that causal structure [Wolfe2016inflation] (see Refs. [pearl2009causality, fritz2012beyond, Himbeeck2018instrumental] for examples). In Appendix C, we show that the noncontextuality inequality of Eq. 12 for the $\rm PBR$$^{+}$ scenario (Fig. 1) arises as a classical causal compatibility inequality within the causal structure known as the bilocality scenario [bilocality], depicted in Fig. 4. (Recall that from Eq. (7), this inequality can be expressed in terms of the conditional probabilities $\mathbb{P}(X_{A}X_{B}Y|S_{A}S_{B}T)$ that are native to the bilocality scenario.) That is, if one is certain that the causal structure of one’s experiment is as shown in Fig. 4, then one does not even need to appeal to noncontextuality to derive the bound in Eq. (12)—rather, the bound follows from the causal structure itself.
@@
One way to obtain a nontrivial classical causal compatibility inequality (i.e., one that can be violated by quantum theory) in the bilocality scenario is to piggyback on Bell’s theorem: by postselecting on the joint measurement’s outcome, one can achieve entanglement swapping, effectively reducing the scenario to a standard Bell scenario with entanglement between $A^{\prime}$ and $B^{\prime}$. It follows that one can achieve every variety of proof that is achievable in the standard Bell scenario, including possibilistic proofs (the paradigm example of which is Hardy’s proof [Hardy1993paradox]) wherein the quantum-classical gap is manifest already at the level of the support of the distribution, i.e., a specification of which outcomes are possible and which are impossible, without regard to the values of the nonzero probabilities. However, the question of interest has always been whether there are novel proofs of quantum-classical gaps in the bilocality scenario, i.e., proofs which do not piggyback on Bell’s theorem through the entanglement swapping protocol [bilocality, BilocalCorrelations, ciudadalanon2024escapingshadowbellstheorem]. The quantum-classical gap we have manifested here is of this variety. Furthermore, to our knowledge, it is the first example of a novel proof of nonclassicality in the bilocality scenario that is possibilistic.
@@
Acknowledgements— We thank Kzysztof Sienicki for pointing out an calculational error and some typos in an earlier version of the paper. DS, RWS, and YY were supported by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. YY was also supported by the Natural Sciences and Engineering Research Council of Canada (Grant No. RGPIN-2024-04419).
@@ Appendix A Overlapping ontic supports
As proven in Ref. [schmidContextual2018], $\Lambda_{0}\neq\Lambda_{+}$, thus, there exists $\lambda^{\ast}\in\Lambda_{0}$ such that $\lambda^{\ast}\notin\Lambda_{+}$, i.e., $\mu_{0}(\lambda^{\ast})>0$ while $\mu_{+}(\lambda^{\ast})=0$. Then, together with the fact that $\mu_{1}(\lambda)$ is nonnegative, Eq. 11 implies that $\mu_{-}(\lambda^{\ast})>0$. Thus, $\lambda^{\ast}$ is in the ontic support for both $\mu_{0}$ and $\mu_{-}$, implying that $\Omega_{0-}$ has nontrivial ontic overlap.
@@ Appendix B Proof of the noncontextuality inequality
Here we derive the noncontextuality inequality on CE, namely, Eq. 12, and show that it is tight, meaning that the maximal value of CE that can be achieved for any theory that admits of a noncontextual ontological model is 3.75. Such a proof does not assume the correctness of quantum theory, so we first translate the scenario described in the main text to one for a general GPT.
@@
$\frac{1}{2}\vec{s}_{0}+\frac{1}{2}\vec{s}_{1}=\frac{1}{2}\vec{s}_{+}+\frac{1}{2}\vec{s}_{-}.$ (13)
@@
${\rm CE}_{T}\coloneqq 1-\frac{1}{4}\sum^{4}_{Y=1}\mathbb{P}(\vec{e}_{Y|T}|\vec{s}_{Y|T})=1-\frac{1}{4}\sum^{4}_{Y=1}\vec{e}_{Y|T}\cdot\vec{s}_{Y|T}.$ (14)
@@
The ontic state space of a bipartite system is simply the Cartesian product, namely, $\Lambda_{A}\times\Lambda_{B}=\Lambda\times\Lambda$, and the ontic representation of a product state is the product of the ontic representations of the unipartite state in the product. We denote the response function representing the measurement effect $\vec{e}_{Y|T}$ as $\xi_{Y|T}(\lambda_{A},\lambda_{B})$ and denote the probability distribution representing the GPT state on $AB$, $\vec{s}_{Y|T}$, as $\mu_{Y|T}(\lambda_{A},\lambda_{B})$, such that
@@
${\rm CE}_{0+}\leq$ $1-\frac{1}{4}\int_{\Lambda_{0}\cap\Lambda_{+}}\int_{\Lambda_{0}\cap\Lambda_{+}}[\xi_{1|0+}(\lambda_{A},\lambda_{B})\mu_{0}(\lambda_{A})\mu_{0}(\lambda_{B})+\xi_{2|0+}(\lambda_{A},\lambda_{B})\mu_{0}(\lambda_{A})\mu_{+}(\lambda_{B})$
$+\xi_{3|0+}(\lambda_{A},\lambda_{B})\mu_{+}(\lambda_{A})\mu_{0}(\lambda_{B})+\xi_{4|0+}(\lambda_{A},\lambda_{B})\mu_{+}(\lambda_{A})\mu_{+}(\lambda_{B})]d\lambda_{A}d\lambda_{B}$
$\leq$ $1-\frac{1}{4}\int_{\Lambda_{0}\cap\Lambda_{+}}\int_{\Lambda_{0}\cap\Lambda_{+}}\sum^{4}_{Y=1}\xi_{Y|0+}(\lambda_{A},\lambda_{B})\min\{\mu_{0}(\lambda_{A}),\mu_{+}(\lambda_{A})\}\min\{\mu_{0}(\lambda_{B}),\mu_{+}(\lambda_{B})\}d\lambda_{A}d\lambda_{B}$
$=$ $1-\frac{1}{4}\left[\int_{\Lambda_{0}\cap\Lambda_{+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2}.$ (20)
@@
${\rm CE}_{0-}\leq 1-\frac{1}{4}\left[\int_{\Lambda_{0}\cap\Lambda_{-}}\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2},$ (22)
${\rm CE}_{1+}\leq 1-\frac{1}{4}\left[\int_{\Lambda_{1}\cap\Lambda_{+}}\min\{\mu_{1}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2},$ (23)
${\rm CE}_{1-}\leq 1-\frac{1}{4}\left[\int_{\Lambda_{1}\cap\Lambda_{-}}\min\{\mu_{1}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2}.$ (24)
@@
${\rm CE}\leq 4-\frac{1}{4}\left\{\left[\int_{\Lambda_{0}\cap\Lambda_{+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2}+\left[\int_{\Lambda_{0}\cap\Lambda_{-}}\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2}\right$.
$\left.+\left[\int_{\Lambda_{1}\cap\Lambda_{+}}\min\{\mu_{1}(\lambda),\mu_{+}(\lambda)\}d\lambda\right]^{2}+\left[\int_{\Lambda_{1}\cap\Lambda_{-}}\min\{\mu_{1}(\lambda),\mu_{-}(\lambda)\}d\lambda\right]^{2}\right\}.$ (25)
@@
$\beta_{0+}\coloneqq\int_{\Lambda_{0}\cap\Lambda_{+}}\min\{\mu_{0}(\lambda),\mu_{+}(\lambda)\}d\lambda,\quad\beta_{0-}\coloneqq\int_{\Lambda_{0}\cap\Lambda_{-}}\min\{\mu_{0}(\lambda),\mu_{-}(\lambda)\}d\lambda,$ (26)
$\beta_{1+}\coloneqq\int_{\Lambda_{1}\cap\Lambda_{+}}\min\{\mu_{1}(\lambda),\mu_{+}(\lambda)\}d\lambda,\quad\beta_{1-}\coloneqq\int_{\Lambda_{1}\cap\Lambda_{-}}\min\{\mu_{1}(\lambda),\mu_{-}(\lambda)\}d\lambda.$ (27)
@@
$\big(\beta_{0+}^{2}+\beta_{0-}^{2}+\beta_{1+}^{2}+\beta_{1-}^{2}\big)\big(1^{2}+1^{2}+1^{2}+1^{2}\big)\geq\big(\beta_{0+}+\beta_{0-}+\beta_{1+}+\beta_{1-}\big)^{2}.$ (29)
@@
${\rm CE}\leq 4-\frac{1}{16}\big(\beta_{0+}+\beta_{0-}+\beta_{1+}+\beta_{1-}\big)^{2}.$ (30)
@@
$\beta_{0+}=$ $\int_{\Lambda_{0+}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{0+-}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+}}\min\{\mu_{0},\mu_{+}\}d\lambda+\int_{\Lambda_{01+-}}\min\{\mu_{0},\mu_{+}\}d\lambda.$ (31)
@@
$\Lambda_{1+}\coloneqq\Lambda_{1}\cap\Lambda_{+}\setminus(\Lambda_{01+-}\cup\Lambda_{01+}\cup\Lambda_{1+-}),$ (35)
$\Lambda_{1-}\coloneqq\Lambda_{1}\cap\Lambda_{-}\setminus(\Lambda_{01+-}\cup\Lambda_{01-}\cup\Lambda_{1+-}),$ (36)